### v1

Abstract
In this paper we introduce the $q$-differential operator for $q$-spinor variables. We establish the $q$-spinor chain rule , the new $q$-differential operator, the $q$-Dirac differential operators and the $q$-complex spinor integrals. We also define the $q$-spinor differential equation. The suggestions for further work at the end of the paper.
Keywords:
$q$-Differential operators, $q$-Dirac operator, $q$-spinor variables, integral formulas in $q$-spinor variables, differential equation in $q$-spinor variables.
msclass: 81Q99, 46E99, 35A24, 15A66 , 16T99 . 17B37.
1 Introduction
Based on the work of Beretetskii et al., Lachieze-Rey, Gori et al., and Cartan, the spinor $\psi^{\alpha}$ is defined as a magnitude components $\alpha=1,2$ expressed as $\psi^{\alpha}=\begin{bmatrix}\psi_{1}\\ \psi_{2}\end{bmatrix}$ and its complex conjugate $\dot{\varphi}$ in terms of the rotation matrices (see [2],[1],[9],[5] for more details). Based on the work of the previously mentioned authors, there are two types of operations on spinors, which are reflections and rotations. In group theory, the set of rotations described by the matrices with complex entries is group $SU(2)$, whose generators are the Pauli matrices, described in the work of Zettili [17]. With respect to rotation matrices, Gori et al. mention, in their work, the rotation matrices that originated the Pauli matrices in the form:
$R_{x}(\theta)=\begin{bmatrix}\cos(\theta/2)&i\sin(\theta/2)\\ i\sin(\theta/2)&\cos(\theta/2)\end{bmatrix},R_{y}(\theta)=\begin{bmatrix}\cos(\theta/2)&\sin(\theta/2)\\ \sin(\theta/2)&\cos(\theta/2)\end{bmatrix},R_{z}(\theta)=\begin{bmatrix}e^{i\theta/2}&0\\ 0&e^{-i\theta/2}\end{bmatrix},$ (1)
being $\theta$ the angle of rotation [5]. Beretetskii et al. define the covariance and contravariance over the spinors by the relation $\psi^{\prime 1}=\psi_{2},\psi^{\prime 2}=-\psi_{1}$ from the matrix $g_{\alpha\beta}=\begin{bmatrix}0&1\\ -1&0\end{bmatrix}$ and, similarly, for pointed spinors, $\psi_{\dot{1}}=\psi^{\dot{2}},\psi_{\dot{2}}=-\psi^{\dot{1}}$ [1]. The same author defines bispinors as the pair $(\psi^{\alpha},\varphi_{\dot{\alpha}})$, which form a broader group of Lorentz, and, with them, the scalar product is formulated as $(\psi^{\alpha},\varphi_{\dot{\alpha}})\cdot(f^{\alpha},h_{\dot{\alpha}})$. The author of the reference [9] mentions algebra $Cl(3)$ as a space-time formulation generated by vectors $e_{\mu}$, which form a basis for $\mathds{R}^{1,3}$ that satisfies the relation $e_{\mu}\cdot e_{\nu}=g_{\mu\nu}$, inducing a 16 -membered basis, as described below:
  • 1
    $1$ scalar,
  • 2
    $(e_{0},e_{1},e_{2},e_{3})$ 4-vector,
  • 3
    $(e_{0}e_{1},e_{0}e_{2},e_{0}e_{3},e_{1}e_{2},e_{2}e_{3},e_{3}e_{1})$ 6- bivectors,
  • 4
    $(e_{1}e_{2}e_{3},e_{0}e_{2}e_{3},e_{0}e_{1}e_{3},e_{0}e_{1}e_{2})$ 4-trivectors,
  • 5
    $e_{5}\equiv e_{0}e_{1}e_{2}e_{3}$ pseudoescalar.
In accordance with the above, the same author describes the Weyl spinors as
$\begin{bmatrix}\psi^{1}\\ \psi^{2}\end{bmatrix}$ and the Dirac spinor $\begin{bmatrix}\psi^{\alpha}\\ \varphi^{\dot{\alpha}}\end{bmatrix}$, as defined in the work of Beretetskii et al [1].
The $q$ Lorentzian algebra was defined in the reference [12]. The quantum complex spinors have components $\psi_{1}$ and $\psi^{2}$ and conjugates $\varphi^{\dot{1}}$ and $\varphi_{\dot{2}}$. For all $q\in\mathds{R}-\left\{0\right\}$, they satisfy the following $q$ relations
$\psi_{1}\psi^{2}=q\psi^{2}\psi_{1},$ $\ \psi^{2}\varphi^{\dot{1}}=q\varphi^{\dot{1}}\psi^{2},$ (2)
$\psi_{1}\varphi^{\dot{1}}=\varphi^{\dot{1}}\psi_{1}-q(q+q^{-1})^{1/2}\varphi_{\dot{2}}\varphi^{\dot{1}},$ $\ \psi^{2}\varphi^{\dot{1}}=\varphi^{\dot{1}}\psi^{2},$ (3)
$\psi_{1}\varphi_{\dot{2}}=q\varphi_{\dot{2}}\psi_{1},$ $\ \varphi^{\dot{1}}\varphi_{\dot{2}}=q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}}.$ (4)
Considering spinors $\psi^{\alpha},\varphi^{\dot{\alpha}}$ and $\left\{\tau^{1},T^{2},S^{1},\sigma^{2}\right\}$ as the generators of the $q$ Lorentzian algebra for the group $U_{q}(su(2))$ [13], we have:
  • 1.
    For $\psi^{\alpha},\alpha=1,2$
    $\tau^{1}\psi_{1}=$ $\ \psi_{1}\tau^{1},$ (5)
    $\tau^{1}\psi^{2}=$ $\ \psi^{2}\tau^{1}-q(q+q^{-1})^{2}\psi_{1}T^{2},$ (6)
    $T^{2}\psi_{1}=$ $\ q^{-1}\psi_{1}T^{2},$ (7)
    $S^{1}\psi_{1}=$ $\ q\psi_{1}S^{1},$ (8)
    $T^{2}\psi^{2}=$ $\ q\psi^{2}T^{2},$ (9)
    $S^{1}\psi^{2}=$ $\ q^{-1}\psi^{2}S^{1}-\psi_{1}\sigma^{2},$ (10)
    $\sigma^{2}\psi_{1}=$ $\ \psi_{1}\sigma^{2},$ (11)
    $\sigma^{2}\psi^{2}=$ $\ \psi^{2}\sigma^{2},$ (12)
  • 2.
    Their complex conjugates $\varphi^{\dot{\alpha}},\alpha=\dot{1},\dot{2}$
    $\tau^{1}\varphi^{\dot{1}}=$ $\ q^{-1}\varphi^{\dot{1}}\tau^{1},$ (13)
    $\tau^{1}\varphi_{\dot{2}}=$ $\ q\varphi_{\dot{2}}\tau^{1},$ (14)
    $T^{2}\varphi^{\dot{1}}=$ $\ \varphi^{\dot{1}}T^{2}+q^{-1}\varphi_{\dot{2}}\tau^{1},$ (15)
    $T^{2}\varphi_{\dot{2}}=$ $\ \varphi_{\dot{2}}T^{2},$ (16)
    $S^{1}\varphi^{\dot{1}}=$ $\ \varphi^{\dot{1}}S^{1},$ (17)
    $\sigma^{2}\varphi^{\dot{1}}=$ $\ q\varphi^{\dot{1}}\sigma^{2}+(q+q^{-1})^{2}\varphi_{\dot{2}}S^{1},$ (18)
    $\sigma^{2}\varphi_{\dot{2}}=$ $\ q\varphi_{\dot{2}}\sigma^{2}.$ (19)
Deformed commutation relations for $q$- Lorentzian algebra are defined in the next proposition, on the quantum-symmetric plane and the quantum anti-symmetric plane.
Proposition 1.1.
Consider generator $T$ of the set $\left\{\tau^{1},T^{2},S^{1},\sigma^{2}\right\}$ for the algebra $U_{q}(su(2))$ and the relations 2, 3, and 4 , defined in [12] and [13] . The $q$ Lorentzian algebra for spinors in the deformed space is defined through the following relations:
$T(\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})=$ $\ (\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})T,$ (20)
$T(\psi_{1}\psi^{2}-q\psi^{2}\psi_{1})=$ $\ (\psi_{1}\psi^{2}-q\psi{2}\psi^{1})T,$ (21)
$T(\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})=$ $\ (\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})T,$ (22)
$T(\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})=$ $\ (\varphi^{\dot{1}}\varphi_{\dot{2}}+q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}})T.$ (23)
Definition 1.2.
The following are the bosonic $q$ deformed Minkowskian Pauli spin matrices defined in the Schmidt work [13]:
$(\sigma^{+})_{\alpha\dot{\beta}}=\begin{bmatrix}0&0\\ 0&q\end{bmatrix},\quad(\sigma^{-})_{\alpha\dot{\beta}}=\begin{bmatrix}q&0\\ 0&0\end{bmatrix},$ (24)
$(\sigma^{3})_{\alpha\dot{\beta}}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\sigma^{0})_{\alpha\dot{\beta}}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (25)
Likewise, the conjugated Pauli matrices are:
$(\overline{\sigma}^{+})_{\dot{\alpha}\beta}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\overline{\sigma}^{-})_{\dot{\alpha}\beta}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (26)
$(\overline{\sigma}^{3})_{\dot{\alpha}\beta}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\overline{\sigma}^{0})_{\dot{\alpha}\beta}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{-1/2}\\ -q^{1/2}&0\end{bmatrix}.$ (27)
The inverse Pauli matrices
$(\sigma^{-1}_{+})_{\alpha\dot{\beta}}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\sigma^{-1}_{-})_{\alpha\dot{\beta}}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (28)
$(\sigma^{-1}_{3})_{\alpha\dot{\beta}}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\sigma^{-1}_{0})_{\alpha\dot{\beta}}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (29)
Finally
$(\overline{\sigma}^{-1}_{+})_{\dot{\alpha}\beta}=\begin{bmatrix}0&0\\ 0&q^{-1}\end{bmatrix},\quad(\overline{\sigma}^{-1}_{-})_{\dot{\alpha}\beta}=\begin{bmatrix}q^{-1}&0\\ 0&0\end{bmatrix},$ (30)
$(\overline{\sigma}^{-1}_{3})_{\dot{\alpha}\beta}=q(q+q^{-1})^{-1/2}\begin{bmatrix}0&q^{1/2}\\ q^{-1/2}&0\end{bmatrix},\quad(\overline{\sigma}^{-1}_{0})_{\dot{\alpha}\beta}=(q+q^{-1})^{-1/2}\begin{bmatrix}0&-q^{-1/2}\\ q^{1/2}&0\end{bmatrix}.$ (31)
Definition 1.3.
The $q$ Lorentzian spinor variables or $q$ spinor variables are defined according to the expressions (2), (3), and (4) as follows:
$u^{2}_{1}\equiv$ $\ \psi_{1}\psi^{2}-q\psi^{2}\psi_{1},$ (32)
$v^{\dot{1}2}\equiv$ $\ \psi^{2}\varphi^{\dot{1}}-q\varphi^{\dot{1}}\psi^{2},$ (33)
$x^{\dot{1}}_{1\dot{2}}\equiv$ $\ \psi_{1}\varphi^{\dot{1}}-\varphi^{\dot{1}}\psi_{1}+q(q+1)^{-1/2}\varphi_{\dot{2}}\varphi^{\dot{1}},$ (34)
$y^{2\dot{1}}\equiv$ $\ \psi^{2}\varphi^{\dot{1}}-\varphi^{\dot{1}}\psi^{2},$ (35)
$z^{\dot{1}}_{\dot{2}}\equiv$ $\ \psi_{1}\varphi_{\dot{2}}-q^{-1}\varphi_{\dot{2}}\varphi^{\dot{1}},$ (36)
$t_{1\dot{2}}\equiv$ $\ \psi_{1}\varphi_{\dot{2}}-q\varphi_{\dot{2}}\psi_{1}.$ (37)
Definition 1.4.
We consider the set $U=\left\{u^{2}_{1},v^{\dot{1}2},x^{\dot{1}}_{1\dot{2}},z^{\dot{1}}_{\dot{2}},y^{2\dot{1}},t_{1\dot{2}}\right\}\subset\mathds{C}$. A function on the $q$ spinor variables is defined as $\Psi(U)=\Psi(u^{2}_{1},v^{\dot{1}2},x^{\dot{1}}_{1\dot{2}},z^{\dot{1}}_{\dot{2}},y^{2\dot{1}},t_{1\dot{2}})$.
Definition 1.5.
Let $f,g:U\longrightarrow\mathds{C}$ be functions and $u^{\beta}\in U$. The following properties are satisfy on the $q$ spinor variables, we state some clear properties of the functions on the $q$ spinor variables
  • 1.
    $(f+g)(u^{\beta})=f(u^{\beta})+g(u^{\beta})$.
  • 2.
    $(f\cdot g)(u^{\beta})=f(u^{\beta})\cdot g(u^{\beta})$.
  • 3.
    $(f-g)(u^{\beta})=f(u^{\beta})-g(u^{\beta})$.
  • 4.
    $\left(\dfrac{f}{g}\right)(u^{\beta})=\dfrac{f(u^{\beta})}{g(u^{\beta})},\quad g(u^{\beta})\neq 0$.
Definition 1.6.
For a function $f:U\longrightarrow\mathds{C}$ and $u^{\beta}\in\mathbb{C}$ , the $q$ spinor derivative is defined as [8] :
$\dfrac{\mathrm{d}_{q}f}{\mathrm{d}_{q}u^{\beta}}=\dfrac{f((qu)^{\beta})-qf(u^{\beta})}{(qu)^{\beta}-qu^{\beta}},$ (38)
and its conjugate complex
$\dfrac{\mathrm{d}_{q}f}{\mathrm{d}_{q}v^{\dot{\alpha}}}=\dfrac{f((qv)^{\dot{\alpha}})-qf(v^{\dot{\alpha}})}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}}.$ (39)
1.1 Clifford algebra and Dirac operator
Let $\left\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n},\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The Clifford algebra is generated over $\mathbb{R}^{n}$ under the relation
$\gamma_{\mu}\gamma_{\nu}+\gamma_{\nu}\gamma_{\mu}=-2\delta_{\mu\nu}\gamma_{0},\quad\gamma_{\mu}^{2}=-|\gamma_{\mu}|^{2}\gamma_{0},\quad\mu,\nu=1,2,...,n,$ (40)
where $\delta_{\mu\nu}$ is the Kronecker symbol (see [11], [3], [10] for more details). We will denote the Clifford algebra by $Cl_{n}$, and each element in $Cl_{n}$ can be expressed by its components as $\sum_{a}\gamma_{a}x_{a}$, where $a=(\mu_{1},...,\mu_{n})$ with each $\mu_{l}\in\left\{1,2,...,n\right\}$. Any element $\textit{{x}}\in\mathds{R}^{n}$ can be identified with a 1-vector in the Clifford algebra [10]
$(x_{1},x_{2},...,x_{n})\longrightarrow\textit{{x}}=x_{1}\gamma_{1}+x_{2}\gamma_{2}+\cdots+x_{n}\gamma_{n}.$ (41)
On other hand, the Dirac operator used here is
$D:=\sum\limits_{\mu=1}^{n}\gamma_{\mu}\frac{\partial}{\partial x_{\mu}},$ (42)
we refer to reader to [3], [6], [4] for more details.
1.2 $q$-Deformed Dirac matrices
Definition 1.7.
The $q$-deformed Dirac matrices are defined in [15], and are given by
$\gamma_{\mu}:=\ \begin{bmatrix}0&(\sigma_{\mu})^{\alpha}_{\dot{\beta}}\\ (\overline{\sigma}_{\mu})^{\dot{\alpha}}_{\beta}&0\end{bmatrix},$ (43)
where $(\sigma_{\mu})^{\alpha}_{\dot{\beta}}$ and $(\overline{\sigma}_{\mu})^{\dot{\alpha}}_{\beta}$ denote the Pauli matrices of $q-$ deformed Minkowski space (e.g. [14] for more details), and are defined as
$\begin{split}(\sigma_{+})^{\alpha}_{\dot{\beta}}=\begin{bmatrix}0&0\\ 0&kq^{1/2}\lambda^{1/2}_{+}\end{bmatrix},\quad(\sigma_{3})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}0&q\\ 1&0\end{bmatrix},\\ (\sigma_{-})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}q^{1/2}\lambda^{1/2}_{+}&0\\ 0&0\end{bmatrix},\quad(\sigma_{0})^{\alpha}_{\dot{\beta}}=k\begin{bmatrix}0&-q^{-1}\\ 1&0\end{bmatrix},\end{split}$ (44)
and their conjugated counterparts
$\begin{split}(\overline{\sigma}_{+})^{\alpha}_{\dot{\beta}}=\begin{bmatrix}0&0\\ 0&\overline{k}q^{-1/2}\lambda^{1/2}_{+}\end{bmatrix},\quad(\overline{\sigma}_{3})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}0&1\\ q^{-1}&0\end{bmatrix},\\ (\overline{\sigma}_{-})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}q^{-1/2}\lambda^{1/2}_{+}&0\\ 0&0\end{bmatrix},\quad(\overline{\sigma}_{0})^{\alpha}_{\dot{\beta}}=\overline{k}\begin{bmatrix}0&1\\ -q&0\end{bmatrix},\end{split}$ (45)
where $k,\overline{k}$ are characteristic parameters associated to bosons ($q=+1$) and fermions ($q=-1$).
1.3 $q$-Spinor complex integral formulas
Definition 1.8.
[8] Let $\Gamma_{q}$ be the closed contour of the deformed quantum complex plane, and $u_{0}^{\beta},v^{\dot{\alpha}}\subset\Gamma_{q}$ point spinors contained in the contour. The $q$- spinor complex integral formulas are defined by
$\oint_{\Gamma_{q}}\dfrac{\Psi(u^{\beta})\mathrm{d}_{q}u^{\beta}}{(qu)^{\beta}-qu^{\beta}_{0}}$ $=\ \dfrac{1}{q}\sum_{n=0}^{\infty}\left[(\overline{\sigma}_{\mu})_{\dot{\alpha}\beta}\Psi(u_{0}^{\beta})\right]^{n},$ (46)
$\oint_{\Gamma_{q}}\dfrac{\Psi((qu)^{\beta})\mathrm{d}_{q}u^{\beta}}{(qu)^{\beta}-qu^{\beta}_{0}}$ $=\ \sum_{n=0}^{\infty}\left[(\overline{\sigma}_{\mu})_{\dot{\alpha}\beta}\Psi((qu_{0})^{\beta})\right]^{n},$ (47)
$\oint_{\Gamma_{q}}\dfrac{\Psi(v^{\dot{\alpha}})\mathrm{d}_{q}v^{\dot{\alpha}}}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}_{0}}$ $=\ \dfrac{1}{q}\sum_{m=0}^{\infty}\left[(\sigma_{\mu})_{\alpha\dot{\beta}}\Psi(v_{0}^{\dot{\alpha}})\right]^{m},$ (48)
$\oint_{\Gamma_{q}}\dfrac{\Psi((qv)^{\dot{\alpha}})\mathrm{d}_{q}v^{\dot{\alpha}}}{(qv)^{\dot{\alpha}}-qv^{\dot{\alpha}}_{0}}$ $=\ \sum_{m=0}^{\infty}\left[(\sigma_{\mu})_{\alpha\dot{\beta}}\Psi((qv_{0})^{\dot{\alpha}})\right]^{m}.$ (49)
Motivation
The topic of this article is $q$-differential operators for $q$-spinor variables, the motivation comes from the study of the differential and integral calculus in $q$-spinor variables studied in [8] and some differential operator, in particular the Dirac operator mentioned in references [3], [4], [6]. In accordance with the above, our interest here is to relate the $q$-spinor variables with some differential operators, especially the $q$-Dirac operator, and its implications with differential calculus, integral and differential equations. The main objective of this work is, therefore, to study the $q$-differential operator on the $q$-spinor variables and its corresponding differential and integral calculus, and its differential equations. Solutions to the differential equation in $q$-spinor variables were also found.
This paper is organized as follows. We briefly recall the preliminaries will be used in this paper in Sect.2. The $q$-differential operators for $q$-spinor variables, the $q$-spinor chain rule, the new $q$-differential operator, the $q$-Dirac differential operator, and the integral formulas in $q$-spinor variables are then proposed in Sect. 3. Finally in the last Section the discussion and some suggestions for further work are presented.
Notation
Taking into account the Definition 1.5 the functions on the $q$ spinor variables we will denote by $\Psi(u^{\alpha}_{\dot{\beta}})$ for all $\alpha,\dot{\beta}=1,2$ , and thus we can write (38) and (39) of the form
$\dfrac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}=\dfrac{\Psi((qu)^{\alpha}_{\dot{\beta}})-q\Psi(u^{\alpha}_{\dot{\beta}})}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}},\quad(qu)^{\alpha}_{\dot{\beta}}\neq qu^{\alpha}_{\dot{\beta}}.$ (50)
2 $q$-Differential operators for $q$-spinor variables
The aim of this section is to define the $q$-differential operator for $q$- spinor variables. To begin, first we will mention an important rule of the $q$- spinor differential calculus, that is the q- spinor chain rule.
2.1 The $q$-spinor chain rule
Proposition 2.1.
Let us consider a $q$-spinor function on the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$. The $q$-spinor chain rule can be expressed as
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}.$ (51)
Proof.
Consider the following derivative
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))-q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))}{(qx_{\mu})-q(x_{\mu})},$ (52)
multiplying by $\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}$, we can rewrite (52) as
$\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))-q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qx_{\mu})-q(x_{\mu})},$ (53)
in virtue of (38) and denoting $\frac{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{\mu})}{(qx_{\mu})-q(x_{\mu})}$ by $\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}}$, finally we have
$\frac{\partial^{q}\Psi}{\partial^{q}x_{j}}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{j}},$ (54)
which completes the proof. ∎
Now, we will define a new $q$-differential operator for $q$-spinor variables over an orthonormal basis of $\mathds{R}^{n}$, different to usual Dirac operator and Cauchy - Riemann operator mentioned in [3], [4], [6], and [7].
2.2 The new $q$-differential operator for $q$-spinor variables
The motivation comes from the construction of the any diiferential operator that satisfy the property $D_{q}^{2}=-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$ for all $\mu,\nu=1,2,...,n$ over an orthonormal basis of $\mathds{R}^{n}$.
Proposition 2.2.
Let $\left\{\mathrm{e}_{1},\mathrm{e}_{2},...,\mathrm{e}_{n}\right\}$ be an orthonormal basis of $\mathds{R}^{n}$. The $q$- Differential operator $D^{q}$
$D_{q}=\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}},$ (55)
is defined over $\mathds{R}^{n}$ under the follows relations
$\mathrm{e}^{2}_{\mu}\frac{\partial_{q}^{2}}{\partial_{q}x^{2}_{\nu}}+\mathrm{e}^{2}_{\nu}\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}$ $=\ -2\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}},$ (56)
$\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}$ $=\ \delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},\quad\mu,\nu,\alpha=1,2,...n.$ (57)
Proof.
The proof is based on the following observation: the square of (55) is equivalent to $-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$. Therefore, we first compute $D^{2}_{q}$ resulting
$D^{2}_{q}$ $=\ \mathrm{e}^{2}_{\mu}\frac{\partial_{q}^{2}}{\partial_{q}x^{2}_{\nu}}+\mathrm{e}^{2}_{\nu}\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}}\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}+\mathrm{e}_{\nu}\frac{\partial_{q}}{\partial_{q}x_{\mu}}\mathrm{e}_{\mu}\frac{\partial_{q}}{\partial_{q}x_{\nu}},$ (58)
and substituting (56) and (57) into (58) we get
$D^{2}_{q}$ $=\ -2\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}+\delta_{\mu\alpha}\frac{\partial^{2}_{q}}{\partial_{q}x_{\mu}\partial_{q}x_{\alpha}},$ (59)
and for $\mu=\alpha$ finally we obtain $-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x^{2}_{\nu}}$, which completes the proof. ∎
According to above, we can said that this operator acts over a $q$-spinor function on the form $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))$. Therefore (55) can be expressed as
$D_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))=\mathrm{e}_{\nu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\mu}}+\mathrm{e}_{\mu}\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\nu}}.$ (60)
and the derivatives $\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\mu}}$ and $\frac{\partial_{q}\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu},x_{\nu}))}{\partial_{q}x_{\nu}}$ can be determined using the $q$-spinor chain rule defined by (51). We will consider the following example
Example 2.3.
Let $\Psi(x_{\mu},x_{\nu})=\exp{ix_{\mu}}$ ) and $u^{1}_{\dot{2}}=q^{2}x_{\mu}$. Now, to obtain the differential operator (55), first we write $\Psi(u^{1}_{\dot{2}})$ as $\exp{iu^{1}_{\dot{2}}}$ ). Later we apply (51)
$\frac{\partial_{q}\Psi}{\partial_{q}x_{\mu}}$ $=\ \frac{\partial_{q}}{\partial_{q}u^{1}_{\dot{2}}}(\exp{iu^{1}_{\dot{2}}})\frac{\partial_{q}}{\partial_{q}x_{\mu}}(q^{2}x_{\mu}),$
[…]
$=\ q^{2}\frac{\exp{iqu^{1}_{\dot{2}}}-q\exp{iqu^{1}_{\dot{2}}}}{[qu]^{1}_{\dot{2}}-q(u^{1}_{\dot{2}})},$
finally we apply (55) to obtain
$D_{q}\Psi(u^{1}_{\dot{2}})=\mathrm{e}_{\nu}q^{2}\frac{\exp{iqu^{1}_{\dot{2}}}-q\exp{iqu^{1}_{\dot{2}}}}{[qu]^{1}_{\dot{2}}-q(u^{1}_{\dot{2}})}.$
The expression (55) can be applied to functions that not depends of $u^{\alpha}_{\dot{\beta}}$, That is, it depends only on the variables $x_{\mu},x_{\nu}$. To obtain this result, it is necessary to define the $q$- derivatives in terms of the $q$-spinor variable $x^{\alpha}$ and $x_{\dot{\beta}}$, and the elements $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$ in the following definition
Definition 2.4.
Let us consider the function $\psi(x_{\mu},x_{\nu})$. The expressions:
$\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=\ \frac{\psi(x_{\mu}+q\mathrm{e}_{\mu}x^{\alpha})-\psi(x_{\mu})}{x^{\alpha}},$ (61)
$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\ \frac{\psi(x_{\nu}+q\mathrm{e}_{\nu}x_{\dot{\beta}})-\psi(x_{\nu})}{x_{\dot{\beta}}},$ (62)
are the $q$-derivatives respect to $x_{\mu}$ and $x_{\nu}$ in terms of $q$- spinor variables $x_{\dot{\beta}},x^{\alpha}$ and the elements $\mathrm{e}_{\mu},\mathrm{e}_{\nu}$.
Let us consider the following remark.
Remark 2.5.
Taking into account the above claim, we can said that the differential operator not apply only over the functions $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$, but also over the function $\psi(x_{\mu},x_{\nu})$, and therefore the operator (55) also depends of the variables $x_{\mu}$ and $x_{\nu}$.
Example 2.6.
Consider the following function $\psi(x_{\mu},x_{\nu})=qx_{\nu}x_{\dot{\beta}},\dot{\beta}=\dot{2}$. Now, to determine the derivatives, we apply (61) and (62)
$\frac{\partial_{q}\psi}{\partial_{q}x_{\mu}}$ $=0$
$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\ \frac{q(x_{\nu}+q\mathrm{e}_{\nu}x_{\dot{2}})x_{\dot{2}}-qx_{\nu}x_{\dot{2}}}{x_{\dot{2}}}$
$=\ q^{2}\mathrm{e}_{\nu}x_{\dot{2}},$
thus the operator (55) is expressed as
$D_{q}\psi=q^{2}x_{\dot{2}}\mathrm{e}_{\mu}\mathrm{e}_{\nu}.$
2.3 The $q$-Dirac differential operator
Definition 2.7.
The $q$-analogue of (42) can be expressed as
$D_{\mu}^{q}=\gamma_{\mu}\frac{\partial^{q}}{\partial^{q}x_{\mu}}.$ (63)
We can now formulate our main results in the following propositions.
2.4 The $q$-differential operators for $q$-spinor variables
Proposition 2.8.
Let $\Psi$ be a function on the $q$-spinor variables. The operator (55) for $q$-spinor variables $D^{q}\Psi$ is given by
$D^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}u^{\alpha}_{\dot{\beta}}.$ (64)
Proof.
Let us consider the expressions (55) and
$\frac{\partial^{q}\Psi}{\partial^{q}x}=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}.$ (65)
Multiplying the left-hand side by $\mathrm{e}_{\nu}$ in (54), we have
$\begin{split}\mathrm{e}_{\nu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}&=\ \mathrm{e}_{\nu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\mathrm{e}_{\nu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}},\end{split}$ (66)
and multiplying the left-hand side by $\mathrm{e}_{\mu}$ in (65) we get
$\begin{split}\mathrm{e}_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\nu}}&=\ \mathrm{e}_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\mathrm{e}_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}},\end{split}$ (67)
adding (66) with (67) and considering (55), finally results
$D^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}u^{\alpha}_{\dot{\beta}},$
which our assertion. ∎
Remark 2.9.
The above proof implies the following relation
$\mathrm{e}_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}-\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\nu}}\mathrm{e}_{\mu}=0.$ (68)
Remark 2.10.
If $\Psi$ also depends of $x_{\nu}$, then
$D^{q}\Psi=\frac{\partial_{q}\Psi}{\partial^{q}x_{\nu}}D^{q}x_{\nu}.$ (69)
Remark 2.11.
The $q$-differential for the coordinate $x_{\mu}$ is given by
$D^{q}x_{\nu}:=\mathrm{e}_{\mu}\mathrm{d}^{q}x_{\nu},$ (70)
consequently the $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
$D^{q}u^{\alpha}_{\dot{\beta}}:=\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\nu}}D^{q}x_{\nu}.$ (71)
Proposition 2.12.
Let $\Psi$ be a function on the $q$-spinor variables. The Dirac operator for $q$-spinor variables $D_{\mu}^{q}\Psi$ is given by
$D^{q}_{\mu}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}.$ (72)
Proof.
Multiplying the left-hand side by $\gamma_{\mu}$ in (54), we have
$\begin{split}\gamma_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}x_{\mu}}&=\ \gamma_{\mu}\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\\ &=\ \frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\gamma_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}},\end{split}$ (73)
and considering (63), finally we get
$D_{\mu}^{q}\Psi=\frac{\partial^{q}\Psi}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}_{\mu}u^{\alpha}_{\dot{\beta}},$
which is our claim. ∎
Remark 2.13.
The above proof implies the following relation
$\gamma_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}-\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial_{q}x_{\mu}}\gamma_{\mu}=0.$ (74)
Remark 2.14.
The Dirac $q$-differential for the coordinate $x_{\mu}$ is given by
$D^{q}_{\mu}x:=\gamma_{\mu}\mathrm{d}^{q}x^{\mu},$ (75)
consequently the Dirac $q$-differential for $u^{\alpha}_{\dot{\beta}}$ is defined as
$D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}:=\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x^{\mu}}D^{q}_{\mu}x^{\mu}.$ (76)
Now, by (52), (64) and (72), we have the follows $q$-spinor integral formulas showed in the following subsection.
2.5 The integral formulas in $q$-spinor variables
Proposition 2.15.
Let $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$ be a $q$-spinor function, and let $\Gamma_{q}$ be the closed contour of the deformed quantum complex plane, and $x_{0}\in\Gamma_{q}$. The integral formulas of the $q$-spinor variables are given by
$\oint_{\Gamma_{q}}\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}$ $=\ \sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(qu^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n},$ (77)
$\oint_{\Gamma_{q}}\frac{\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{qu^{\alpha}_{\dot{\beta}}(x_{\mu})-(qu)^{\alpha}_{\dot{\beta}}(x_{0})}$ $=\ \frac{1}{q}\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n},$ (78)
where $\gamma_{\mu}$ are the $q$-deformed Dirac matrices defined in (43).
Proof.
The proof of this result is based on the ideas presented in [8]. Combining (38) with (72) we get
$D^{q}_{\mu}\Psi=\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}}-\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}-qu^{\alpha}_{\dot{\beta}}},$ (79)
now, we integrate over the closed contour $\Gamma_{q}$ and we take $x_{0}\in\Gamma_{q}$ to obtain
$\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi=\oint_{\Gamma_{q}}\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}-\oint_{\Gamma_{q}}\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}.$ (80)
Hence, to solve the integral $\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8], interchanging the Pauli matrices of $q$-deformed Minkowski space (e.g. [14]) by the $q$-deformed Dirac matrices (43), obtaining
$\begin{split}\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}-\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}&=\\ \oint_{\Gamma_{q}}\frac{\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})}-\oint_{\Gamma_{q}}\frac{q\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}}{(qu)^{\alpha}_{\dot{\beta}}(x_{\mu})-qu^{\alpha}_{\dot{\beta}}(x_{0})},\end{split}$ (81)
and finally, equalating terms that depend on $\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{\mu})$ and $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$ we obtain (77) and (78), and the proof is complete. ∎
We will mention an important consequence of the Definition 3.4. starting of Eqs. (77) and (78): the formulation of the integral $\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}}(x))d^{q}x_{\mu}$, which we will mention in the following theorem.
Theorem 2.16.
Let $\Gamma_{q}$ be a closed contour and suppose that $x_{0}\in\Gamma_{q}$. Then for a function on $q-$ spinor variables $\Psi(u^{\alpha}_{\dot{\beta}}(x))$ the $q$-spinor integral formula is given by
$\frac{1}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},\quad b\neq 0.$ (82)
Proof.
To formulate we can consider the following differential equation in $q$- spinor variables
$D^{q}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}})-b\Psi(u^{\alpha}_{\dot{\beta}})=0.$ (83)
Multiplying on both sides of (83) by $\mathrm{d}^{q}x_{\mu}$, and using (63) we get
$\begin{split}\frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D^{q}_{\mu}u^{\alpha}_{\dot{\beta}}\mathrm{d}^{q}x_{\mu}&=\ b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu}\\ \frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}\gamma_{\mu}\frac{\partial^{q}u^{\alpha}_{\dot{\beta}}}{\partial^{q}x_{\mu}}\mathrm{d}^{q}x_{\mu}&=\ b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},\end{split}$ (84)
taking into account (74), (75) and (76), we can rewrite (84) as
$\frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}D_{\mu}^{q}u^{\alpha}_{\dot{\beta}}=b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},$ (85)
from (38) it follows that
$\frac{\partial^{q}\Psi(u^{\alpha}_{\dot{\beta}})}{\partial^{q}u^{\alpha}_{\dot{\beta}}}=\dfrac{q\Psi(u^{\alpha}_{\dot{\beta}})}{qu^{\alpha}_{\dot{\beta}}-(qu)^{\alpha}_{\dot{\beta}}},$ (86)
substituting (86) into (85) gives
$\left[\frac{q\Psi(u^{\alpha}_{\dot{\beta}})}{qu^{\alpha}_{\dot{\beta}}-(qu)^{\alpha}_{\dot{\beta}}}\right]D_{\mu}^{q}u^{\alpha}_{\dot{\beta}}=b\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},$ (87)
we continue in this fashion integrating over $\Gamma_{q}$ on both sides for $x_{0}\in\Gamma_{q}$ , and using (78), finally we get
$\frac{1}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x_{\mu},\quad b\neq 0,$
which is our claim. This expression is called the $q$-spinor integral formula for $\Psi(u^{\alpha}_{\dot{\beta}}(x_{\mu}))$. ∎
The same reasoning applies to the differential equation $D^{q}_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}})-b\Psi((qu)^{\alpha}_{\dot{\beta}})=0$ to obtain the integral in $q$-spinor variables
$\frac{1}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi((qu)^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi((qu)^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x_{\mu}.$ (88)
Remark 2.17.
The expression (82) also can be expressed in virtue of (75) as
$\frac{\gamma^{\mu}}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}})D^{\mu}_{q}x.$ (89)
Notice that the expression (82) is not implies the final solution of (83).
Remark 2.18.
Similar arguments apply to the new $q$-differential operator (55), resulting
$\frac{\mathrm{e}^{\mu}}{qb}\left\{\sum\limits_{n=0}^{\infty}\left[\mathrm{e}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\oint_{\Gamma_{q}}\Psi(u^{\alpha}_{\dot{\beta}})D^{q}_{\mu}x.$ (90)
Proof.
It is sufficient to replace $\gamma_{\mu}$ by $\mathrm{e}_{\mu}$, to obtain (90), considering the differential equation in $q$-spinor variables of the form
$D^{q}\Psi(u^{\alpha}_{\dot{\beta}})-b\Psi(u^{\alpha}_{\dot{\beta}})=0,$ (91)
where $D^{q}$ is the new $q$-differential operator given by (55). ∎
However we will can solve differential equations in $q$-spinor variables of the form
$D^{q}_{\mu}\psi(u^{\alpha}_{\dot{\beta}})-b\phi(u^{\alpha}_{\dot{\beta}})$ = 0,absent 0=\ 0,= 0 , (92)
$D^{q}\psi(u^{\alpha}_{\dot{\beta}})-a\phi(u^{\alpha}_{\dot{\beta}})$ = 0absent 0=\ 0= 0 (93)
which we will show in the following section.
3 Differential equations in $q$-spinor variables
In order to obtain the solution of (92), it is necessary to put the following condition on $\psi$
$\oint_{\Gamma_{q}}D_{\mu}^{q}\psi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x^{\mu}=\psi(u^{\alpha}_{\dot{\beta}}(x_{0})),\quad x_{0}\in\Gamma_{q}.$ (94)
Therefore, integrating both sides with respect to $x^{\mu}$ in (92) , applying (94), we get
$\begin{split}\oint_{\Gamma_{q}}D^{q}_{\mu}\psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu}&=\ b\oint_{\Gamma_{q}}\phi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu},\\ \psi(u^{\alpha}_{\dot{\beta}}(x_{0}))&=\ b\oint_{\Gamma_{q}}\phi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu},\\ \psi(u^{\alpha}_{\dot{\beta}}(x_{0}))&=\ b\oint_{\Gamma_{q}}\phi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu}\end{split}$ (95)
and applying (82) we obtain
$\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))=\frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\phi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}.$ (96)
Lemma 3.1.
We can generalize the solution (96) for all $x\in\Gamma_{q}$ as
$\psi(u^{\alpha}_{\dot{\beta}}(x))=\frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\phi(u^{\alpha}_{\dot{\beta}}(x))\right]^{n}\right\}.$ (97)
To solve (93) we will consider the following remarks
Remark 3.2.
We first consider the new $q$-differential operator given by (55). Now, to solve the integral $\oint_{\Gamma_{q}}D^{q}\Psi$, we will use similary the proof of the Theorem 2.9 of the reference [8] (also can see the proof of Proposition 3.10.) to obtain similar expressions to (96) and (97), i.e.
$\oint_{\Gamma_{q}}D^{q}\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))$ $=\ \psi(u^{\alpha}_{\dot{\beta}}(x_{0})),$ (98)
$\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))$ $=\ \frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\mathrm{e}_{\mu}\phi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\},$ (99)
Now, we will consider the following examples
Example 3.3.
Consider the differential equation in $q$-spinor variables of the form
$D^{q}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}})-e\gamma^{\mu}A^{q}_{\mu}(x)\Psi(u^{\alpha}_{\dot{\beta}})-mg(u^{\alpha}_{\dot{\beta}})=0,\quad e\in\mathds{R},$ (100)
being $A^{q}_{\mu}(x)$ a $q$-potential function. Now, to solve (100), we can proceed analogously to the solution of (92) applying (94)
$\begin{split}\oint_{\Gamma_{q}}D^{q}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu}-e\gamma^{\mu}\oint_{\Gamma_{q}}A^{q}_{\mu}(x)\Psi(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu}&=\ m\oint_{\Gamma_{q}}g(u^{\alpha}_{\dot{\beta}})\mathrm{d}^{q}x^{\mu},\\ \Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))-e\gamma^{\mu}\oint_{\Gamma_{q}}A^{q}_{\mu}(x)\Psi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x^{\mu}&=\ m\oint_{\Gamma_{q}}g(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x^{\mu},\end{split}$ (101)
and using (82) we obtain finally
$\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))+\frac{e}{qm}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}A^{q}_{\mu}(x_{0})\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}g(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}.$ (102)
Now, let us see other example.
Example 3.4.
Consider the differential equation in $q$-spinor variables (similar to (100)) of the form
$\gamma^{\mu}\partial^{q}_{\mu}\Psi(u^{\alpha}_{\dot{\beta}})-e\gamma^{\mu}A^{q}_{\mu}(x)\Psi(u^{\alpha}_{\dot{\beta}})-m\Psi(u^{\alpha}_{\dot{\beta}})=0,\quad e\in\mathds{R},$ (103)
where $A^{q}_{\mu}(x)$ is the same $q$-potential function of above example. This follows by the same method as in the above example, obtaining
$\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))+\frac{e}{qm}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}A^{q}_{\mu}(x_{0})\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}=\frac{1}{q}\left\{\sum\limits_{n=0}^{\infty}\left[\gamma_{\mu}\Psi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}.$ (104)
Remark 3.5.
The expression (103) can be written as
$(\gamma^{\mu}\mathcal{D}^{q}_{\mu}-m)\Psi(u^{\alpha}_{\dot{\beta}})=0,$ (105)
where $\mathcal{D}^{q}_{\mu}=\partial^{q}_{\mu}-eA^{q}_{\mu}(x)$ is the q- covariant derivative.
Example 3.6.
Let us consider the differential equation in $q$- spinor variables $aD^{q}\psi+b\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))=0$, where $B^{\mu}_{q}(x)$ is some $q$-arbitrary potential function. Therefore
$D^{q}\psi=-\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x)),$ (106)
Using (55) (only the second contribution) we get
$\mathrm{e}_{\mu}\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}$ $=\ -\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x)),$
multiplying both sides by $\mathrm{d}^{q}x_{\nu}$ results and applying (70)
$\mathrm{e}_{\mu}\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}\mathrm{d}^{q}x_{\nu}$ $=\ -\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x_{\nu}$
$\frac{\partial_{q}\psi}{\partial_{q}x_{\nu}}D^{q}x_{\nu}$ $=\ -\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x_{\nu}$
$D^{q}\psi$ $=\ -\frac{b}{a}\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))\mathrm{d}^{q}x_{\nu},$
integrating over the closed contour $\Gamma_{q}$ , considering $x_{0}\subset\Gamma_{q}$ and using (90), (98), (99) and (69) into (106) we get
$\psi(u^{\alpha}_{\dot{\beta}}(x_{0}))=-\frac{1}{aq}\mathrm{e}^{\mu}\left\{\sum\limits_{n=0}^{\infty}\left[\mathrm{e}_{\mu}B^{\mu}_{q}(x_{0})\phi(u^{\alpha}_{\dot{\beta}}(x_{0}))\right]^{n}\right\}.$ (107)
and for all $x\in\Gamma_{q}$
$\psi(u^{\alpha}_{\dot{\beta}}(x))=-\frac{1}{aq}\mathrm{e}^{\mu}\left\{\sum\limits_{n=0}^{\infty}\left[\mathrm{e}_{\mu}B^{\mu}_{q}(x)\phi(u^{\alpha}_{\dot{\beta}}(x))\right]^{n}\right\}.$
4 Discussion and suggestions for further work
In Section 2, the equations (51), (60) and (72) describe some $q$-differential operators for $q$-spinor variables. Respect to (51), we can said that the function on the $q$-spinor variables does not depend on ly on the variable $u^{\alpha}_{\dot{\beta}}$ but also on $x_{\nu}$. From this result, the new $q$- differential operator for $q$-spinor variables expressed by (60) was proposed. This operator is motivated from the construction of the any differential operator that satisfy the property $D_{q}^{2}=-\frac{\partial^{2}}{\partial_{q}x^{2}_{\mu}}-\frac{\partial^{2}_{q}}{\partial_{q}x_{\nu}^{2}}$ for all $\mu,\nu=1,2,...,n$. To obtain $D^{2}_{q}$ it is necessary to use the relations (56) and (57). This operator is different to usual Dirac operator and Cauchy Riemann operator mentioned in [3], [4], [6], and [7]. In the case of the Dirac operator for $q$-spinor variables, defined by (72), has been stablished from the Remarks 2.10 and 2.11. This operator is expressed in terms of the $q$-deformed Dirac matrices mentioned by Schmidt [15]. From the $q$-deformed Dirac operator, we define the integral formulas in $q$-spinor variables with the aim to solve the differential equations in $q$-spinor variables. Physically we can said that the Example 3.4 describes the Dirac equation for the electromagnetic case on the $q$-spinor variables, and furthermore the potential $A^{q}_{\mu}(x)$ can be interpreted as the q- electromagnetic potential. There are two further topics arising from this paper which are worth investigation., there is the problem of describing the Maxwell Electrodynamic Algebra which is defined by the following commutation relations
$A_{\mu}A^{\mu}$ $=\ |A_{0}|^{2}-A_{X}^{2},\quad X=1,2,3,$ (108)
$f_{\mu\nu}$ $=\ \partial_{\mu}A_{\nu}-\partial_{\nu}A_{\mu},$ (109)
$\partial_{\mu}A^{\mu}=\partial_{\nu}A^{\nu}$ = 0,absent 0=\ 0,= 0 , (110)
$D_{\mu}$ $=\ \partial_{\mu}-eA_{\mu},\quad e\in\mathds{R}$ (111)
$\partial^{\mu}f_{\mu\nu}$ $=\ \Gamma_{\nu},$ (112)
$\partial_{\nu}f_{0}^{\mu\nu}$ = 0,absent 0=\ 0,= 0 , (113)
where $f_{0}^{\mu\nu}=\varepsilon^{\mu\nu 0}f_{\mu\nu}$ , $f^{\mu\nu}=0$ if $\mu=\nu$ and $f^{\mu\nu}\neq 0$ in otherwise. Finally, from (103) and taking into account the above, one can propose the $q-$ Dirac - Maxwell algebra, which is subject to relations
$f^{q}_{\mu\nu}$ $=\ \mathcal{D}^{q}_{\mu}A^{q}_{\nu}-\mathcal{D}^{q}_{\nu}A^{q}_{\mu},$ (114)
$\textit{{D}}^{q}$ $=\ \boldsymbol{\partial}^{q}-e\textit{{A}}\quad e\in\mathds{R},$ (115)
being $\textit{{D}}^{q}=\boldsymbol{\gamma}^{\mu}\mathcal{D}^{q}_{\mu},\boldsymbol{\partial}^{q}=\boldsymbol{\gamma}^{\mu}\partial^{q}_{\mu}$ and $\textbf{e}=\gamma^{\mu}A_{\mu}$, where $\boldsymbol{\gamma}^{\mu},\mu=1\cdots n$ are the generators for the Clifford algebras, and (115) is called the Covariant Derivative. Other suggestion is the formulation of the $q-$ Real Spinor Calculus based on the work of Zatloukal [16], which is defined by the following expressions for the derivatives
$\frac{\boldsymbol{\partial}_{q}\psi}{\boldsymbol{\partial}_{q}\textit{{x}}_{\dot{\alpha}}^{\beta}}=\frac{\psi(\textit{{x}}^{\beta}_{\dot{\alpha}}+q\textit{{u}}^{\beta}_{\dot{\alpha}})-q\psi(\textit{{u}}^{\beta}_{\dot{\alpha}})}{\textit{{x}}^{\beta}_{\dot{\alpha}}},$ (116)
where $\textit{{u}}^{\beta}_{\dot{\alpha}}=(\gamma_{\mu}\gamma_{\nu}u)^{\beta}_{\dot{\alpha}}$. The chain rule
$\frac{\boldsymbol{\partial}_{q}\Psi}{\boldsymbol{\partial}_{q}x_{j}}=\frac{\boldsymbol{\partial}_{q}\Psi}{\boldsymbol{\partial}_{q}\textit{{x}}^{\beta}_{\dot{\alpha}}}\frac{\boldsymbol{\partial}_{q}\textit{{x}}^{\beta}_{\dot{\alpha}}}{\boldsymbol{\partial}_{q}x_{j}},$ (117)
the $q$-difference operator for $q-$ real spinor variables
$\textit{{D}}_{2}^{q}$ $=\ \hat{\gamma}_{2}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}x_{2}},$ (118)
$\textit{{D}}^{q}_{j}$ $=\ i\hat{\gamma}_{5}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}x_{j}},$ (119)
$\underline{\textit{{D}}}_{j}^{q}$ $=\ i\hat{\gamma}_{2}\hat{\gamma}_{5}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}x_{j}},\quad j=1,..,5.$ (120)
For a function $\psi:(\textit{{v}}_{\dot{k}},\textit{{p}}_{\dot{k}})\longrightarrow\mathds{R}^{m}$, the $q$-conjugated derivatives are defined as
$\frac{\boldsymbol{\partial}_{q}\psi}{\boldsymbol{\partial}_{q}\textit{{v}}_{\dot{k}}}$ $=\ \frac{\psi(\textit{{v}}_{\dot{k}}+q\textit{{x}}^{\beta}_{\dot{\alpha}})-q\psi(\textit{{x}}^{\beta}_{\dot{\alpha}})}{\textit{{v}}_{\dot{k}}},$ (121)
$\frac{\boldsymbol{\partial}_{q}\psi}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{k}}}$ $=\ \frac{\psi(\textit{{p}}_{\dot{k}}+q\textit{{u}}^{\beta}_{\dot{\alpha}})-q\psi(\textit{{u}}^{\beta}_{\dot{\alpha}})}{\textit{{p}}_{\dot{k}}}.$ (122)
The $q$-difference operators associated to conjugated real spinor variables are given by
$\textit{{D}}_{q}$ $=\ \frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{v}}_{\dot{0}}}+\gamma_{1}\gamma_{3}\frac{\boldsymbol{\partial_{q}}}{\boldsymbol{\partial}_{q}\boldsymbol{v}_{\dot{1}}}+i\gamma_{3}\gamma_{\dot{0}}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{v}}_{\dot{2}}}+\gamma_{1}\gamma_{2}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{v}}_{\dot{3}}},$ (123)
$\textit{{D}}^{\prime}_{q}$ $=\ \frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{0}}}+\gamma_{1}\gamma_{3}\frac{\boldsymbol{\partial_{q}}}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{1}}}+i\gamma_{3}\gamma_{0}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{2}}}+\gamma_{1}\gamma_{2}\frac{\boldsymbol{\partial}_{q}}{\boldsymbol{\partial}_{q}\textit{{p}}_{\dot{3}}},$ (124)
and the $q$-spinor real integral formulas of the $q$-spinor conjugated variables are given by
$\int_{\Omega_{q}}\frac{\psi(q\textit{{v}}_{\dot{k}})\mathbf{d}_{q}\textit{{v}}_{\dot{k}}}{\textit{{v}}_{\dot{k}}-\textit{{x}}^{\beta}_{\dot{\alpha}}}$ $=\ q\sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi(q\textit{{x}}^{\beta}_{\dot{\alpha}})]^{n},$ (125)
$\int_{\Omega_{q}}\frac{\psi[(1-q^{-1})\textit{{v}}_{\dot{k}}]\mathbf{d}_{q}\textit{{v}}_{\dot{k}}}{\textit{{v}}_{\dot{k}}-\textit{{x}}^{\beta}_{\dot{\alpha}}}$ $=\ \sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi[(1-q^{-1})\textit{{x}}^{\beta}_{\dot{\alpha}}]]^{n},$ (126)
$\int_{\Omega_{q}}\frac{\psi(q\textit{{p}}_{\dot{k}})\mathbf{d}_{q}\textit{{p}}_{\dot{k}}}{\textit{{p}}_{\dot{k}}-\textit{{u}}^{\beta}_{\dot{\alpha}}}$ $=\ q\sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi(q\textit{{u}}^{\beta}_{\dot{\alpha}})]^{n},$ (127)
$\int_{\Omega_{q}}\frac{\psi[(1-q^{-1})\textit{{p}}_{\dot{k}}]\mathbf{d}_{q}\textit{{p}}_{\dot{k}}}{\textit{{p}}_{\dot{k}}-\textit{{u}}^{\beta}_{\dot{\alpha}}}$ $=\ \sum\limits_{n=0}^{\infty}[\gamma^{\mu}\gamma^{\nu}\psi[(1-q^{-1})\textit{{u}}^{\beta}_{\dot{\alpha}}]]^{n}.$ (128)